Last Chg Open High Sep 2012 79501547800 7970 Dec 2012 7834 122 7724 7856 Mar 2013 78321147724 7844 May 2013 7796

A) Suppose you buy one contract for March Delivery. If the contract closes in March at a level of 787.25, what will your profit be?

B) How many March Maturity Contracts are outstanding
Answer #1

a). Corn futures list prices in cents per bushel. One contract is of 5,000 bushels.

Last price for March delivery = 783'2 cents (The number after apostrophe represents eighths of a cent) so converting it we get

783 cents + (2*1/8) cents = 783 + 0.25 = 783.25 cents or $7.8325

Closing price is 787.25 cents or $7.8725

Profit = 7.8725 - 7.8325 = 0.04 per bushel

Profit for one contract = profit per bushel*number of bushels = 0.04*5,000 = $200

b). Open interest gives the number of outstanding contracts for each maturity month. For March, it is 135,778.